OpenStudy (anonymous):

Grant can row 6 miles in 36 minutes. Arianna can row 3,520 yards in 20 minutes. In miles per hour, how much faster does Grant row than Arianna?

OpenStudy (anonymous):

1 mile is 1760yds Grant's speed is \[\frac{ 1mile }{ 1\min }\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

3520yd divide by 1760yd you get 2miles

OpenStudy (anonymous):

so arianna's speed is \[\frac{ 1mile }{ 10mins } = \frac{ 0.1mile }{ 1\min }\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

would equal up to be 29.3

OpenStudy (anonymous):

so we convert them mile/min to mph

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

multiple both grant and arianna speed by 60 to get the units to mile per hour. Then: Grant's speed - Arianna's speed

OpenStudy (anonymous):

I got 29.3 for Arianna

OpenStudy (anonymous):

What's your calculation steps?

OpenStudy (anonymous):

5.156.8

OpenStudy (anonymous):

No... Arianna's speed is in yards, you need to convert it to miles

OpenStudy (anonymous):

so convert 29.3 to yards

OpenStudy (anonymous):

I meant mile

OpenStudy (anonymous):

Arianna row 3520 yards in 20mins 1 mile is 1760 yards \[\frac{ 3520yd }{ 1760yd } = 2 miles\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

6195200

OpenStudy (anonymous):

\[\frac{ 2miles }{ 20mins } = \frac{ 1mile }{ 10mins } = \frac{ 0.1miles }{ 1\min }\]

OpenStudy (anonymous):

40

OpenStudy (anonymous):

40?

OpenStudy (anonymous):

I got 40 somehow

OpenStudy (anonymous):

ok we have both Grant and Arianna speed in miles per min (check the working above) now we convert them to miles per hour

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Think I made a typo the top again....at Grant's speed should be \[\frac{ 0.1667miles }{ 1\min }\]

OpenStudy (anonymous):

Multiply Grant's speed by 60 you should get 10miles per hour Multiply Arianna's speed by 60 you should get 6miles per hour

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

10-6=4 which is the answer

OpenStudy (anonymous):

okay